OpenStudy (anonymous):

is the simplest form of: 9 sqrt of (-27x^9) is x of 9sqrt of-27

OpenStudy (anonymous):

like in this picture:

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

you could do more

OpenStudy (anonymous):

put in this form (-27)^(1/9)

OpenStudy (anonymous):

(-3^3)^(1/9) (-3)^(3/9) (-3)^(1/3)

OpenStudy (anonymous):

x 3srq(-3)

OpenStudy (anonymous):

how? @AJ01 like this?

OpenStudy (jhannybean):

\[\large \sf\sqrt[9]{-27x^9} \iff (-27x^9)^{1/9}\]

OpenStudy (jhannybean):

So just distribute the \(\sf 1/9\) power.

OpenStudy (jhannybean):

\[\sf \large (-27)^{1/9} \cdot (x^9)^{1/9}\]\[=\sf [(-3)^3]^{1/9} \cdot x^{9 \cdot 1/9}\]

OpenStudy (anonymous):

\[\sqrt[9]{-27}\] \[\sqrt[9]{-3^{3}}\] \[\sqrt[\frac{ 9}{ 3}]{-3}\] \[\sqrt[3]{-3}\]

OpenStudy (jhannybean):

Following my method: \[\sf [(-3)^3]^{1/9} \cdot x^{9 \cdot 1/9}\]\[\sf =(-3)^{3 \cdot 1/9} \cdot x^1\]\[=(-3)^3 \cdot x\]

OpenStudy (jhannybean):

Just simplify that.